OpenStudy (anonymous):

Change 4x-5y=-20 and x+3y= -3 to y=mx+b form.

OpenStudy (anonymous):

5y=4x+20 y=(4/5)x+4 3y=-x-3 y=(-1/3)x-1

OpenStudy (anonymous):

Why is it positive 4x+20? Wouldn't it be -4x-20?

OpenStudy (anonymous):

2 negatives = positive

OpenStudy (anonymous):

Ohhh okay. Thank you!